### v1

# Can AI-Generated Persuasion Be Detected? Persuaficial Benchmark and AI vs. Human Linguistic Differences

Abstract
Large Language Models (LLMs) can generate highly persuasive text, raising concerns about their misuse for propaganda, manipulation, and other harmful purposes. This leads us to our central question: Is LLM-generated persuasion more difficult to automatically detect than human-written persuasion? To address this, we categorize controllable generation approaches for producing persuasive content with LLMs and introduce Persuaficial, a high-quality multilingual benchmark covering six languages: English, German, Polish, Italian, French and Russian. Using this benchmark, we conduct extensive empirical evaluations comparing human-authored and LLM-generated persuasive texts. We find that although overtly persuasive LLM-generated texts can be easier to detect than human-written ones, subtle LLM-generated persuasion consistently degrades automatic detection performance. Beyond detection performance, we provide the first comprehensive linguistic analysis contrasting human and LLM-generated persuasive texts, offering insights that may guide the development of more interpretable and robust detection tools.
Can AI-Generated Persuasion Be Detected?
Persuaficial Benchmark and AI vs. Human Linguistic Differences
Arkadiusz Modzelewski1,2,3, Paweł Golik, Anna Kołos3, Giovanni Da San Martino1 1University of Padua, Italy 2Polish-Japanese Academy of Information Technology, Poland 3NASK National Research Institute, Poland Correspondence: contact@amodzelewski.com
1 Introduction
Persuasive writing, which uses rhetorical techniques and devices to influence audiences, has become central to modern communication Gass and Seiter (2022). We live in an era where artificial intelligence increasingly shapes propaganda and persuasive communication in news, political discourse, and social media Bergmanis-Korāts et al. (2024); Goldstein et al. (2024). Large Language Models, now widely used in writing and communication tasks, demonstrate a growing potential to produce persuasive text and influence public opinion Pauli et al. (2025); Bai et al. (2025); Breum et al. (2024); Karinshak et al. (2023). Several studies have explored how effectively LLMs can identify persuasive language Sprenkamp et al. (2023); Panasyuk (2025). Yet, to the best of our knowledge, no prior work has addressed whether the automatic detection of LLM-generated persuasion is more challenging than detecting persuasion in human-written texts. Understanding this distinction is crucial, as it reveals the extent to which current detection systems may be vulnerable to increasingly sophisticated AI-driven persuasive content. Furthermore, although previous research has highlighted the importance of mitigating and defending against AI-generated persuasion Burtell and Woodside (2023); El-Sayed et al. (2024), it has largely focused on persuasion detection, leaving the linguistic differences between human-written and LLM-generated persuasive texts unexplored. Understanding these differences could deepen our knowledge of AI-driven persuasion and support the development of more effective automatic detection methods. To address these gaps, we investigate two key research questions: RQ1 Is controllably generated AI persuasion harder for LLMs to detect in a zero-shot setting than human-written persuasion? and RQ2 What are the linguistic differences between controllable LLM-generated and human-written persuasive texts?.
To address our first research question, we introduce Persuaficial, a newly constructed benchmark of artificially generated persuasive texts. Persuaficial is a novel multilingual resource comprising synthetic persuasive content produced using four generation approaches inspired by Chen and Shu (2023). The dataset is created in a controlled manner, leveraging human-written texts drawn from established datasets Piskorski et al. (2023c); Tan et al. (2016); Moral et al. (2023). In our experiments, we evaluate the detectability of AI and human-written persuasive texts using four different LLMs, including commercial closed models and open-weight models. Our analysis focuses on English, but we also provide analysis on five additional languages: German, French, Italian, Polish, and Russian.
To address RQ2, we conducted an analysis using the StyloMetrix tool [^1] https://github.com/ZILiAT-NASK/StyloMetrix, which generates fully interpretable and reproducible vectors representing a wide range of linguistic features in text Okulska et al. (2023). Prior work shows that StyloMetrix performs well on persuasion detection with classical machine learning Modzelewski et al. (2023, 2024), underscoring its suitability for studying the linguistic features of persuasive texts. In our analysis, we examined the full range of linguistic features offered by open-source StyloMetrix.
Our main contributions are summarized as:
  • We introduce Persuaficial, a novel persuasion benchmark with approximately 65k multilingual texts generated via four controllable approaches with four LLMs.
  • We are the first to investigate whether persuasive text generated by LLMs is harder to detect than human-written persuasive text. We conduct this analysis across four LLMs and 16 controllable generation settings, providing a comprehensive evaluation of detection difficulty across diverse persuasive text generation approaches.
  • Our work is the first to characterize the linguistic differences between LLM-generated and human-written English persuasive texts. Our analysis encompasses 196 distinct linguistic features.
We release our codebase, dataset and all prompts222anonymous.4open.science/r/Persuaficial-26EC/.
2 Human Persuasion Datasets
In our analysis, we employed three well-established datasets that had been previously annotated by humans. Using multiple datasets mitigates potential bias that could arise from relying on a single source and ensures a broader coverage of persuasion phenomena. We selected datasets that are widely used and cited in persuasion research. Below, we describe the human-created datasets used in our study:
  • SemEval 2023 Task 3 Dataset: A multilingual, multifaceted collection of online news articles annotated with various persuasion techniques on paragraph level Piskorski et al. (2023d). Its taxonomy and dataset are widely adopted within the NLP community for persuasion research Barrón-Cedeño et al. (2024); Dimitrov et al. (2024); Modzelewski et al. (2025). This dataset was introduced as part of SemEval 2023 Task 3 on persuasion detection Piskorski et al. (2023b).
  • DIPROMATS 2024 Task 1 Dataset: A dataset consisting of posts from the X platform (former Twitter) used for the DIPROMATS 2024 shared task including propaganda detection Moral et al. (2024, 2023). The dataset contains messages from diplomats and authorities of major world powers, including China, the United States, Russia, and the European Union. DIPROMATS 2024 was part of IberLEF, an annual Spanish shared evaluation campaign Chiruzzo et al. (2024).
  • ChangeMyView: A dataset derived from the Reddit ChangeMyView discussion community. Dataset contains 3,051 conversations in which the persuader tries to convince the persuadee to change their mind Tan et al. (2016). It is commonly used in persuasion research Ziems et al. (2024); Wei et al. (2016); Dutta et al. (2020).
Together, these datasets provide diverse perspectives on human persuasion, covering different sources and communicative contexts. Consequently, our analysis captures diverse persuasion understanding, which improves the generalizability of our findings. For further justification of our dataset choices, please refer to Appendix A.
3 Persuaficial: Artificially Generated Persuasion Dataset
In this section, we introduce a new dataset, Persuaficial, which contains persuasion texts generated artificially using large language models.
3.1 Persuasive Text Generation Approaches
Inspired by Chen and Shu (2023), which explores methods for generating synthetic misinformation, we developed a controllable persuasion generation $G_{P}$, in which $P$ denotes one of four approaches for generating persuasive text: Paraphrasing, Rewriting with Subtle Persuasion, Rewriting with Intensified Persuasion, and Open-ended generation.
Each persuasive text was generated under controlled prompting conditions. Specifically, prompts included human-annotated persuasive texts together with instructions to paraphrase or rewrite the input while preserving, strengthening, or softening its persuasive effect. For the open-ended generation, following the approach of Chen and Shu (2023), the model was provided with concise summaries of the corresponding human-annotated persuasive examples and prompted to generate persuasive text freely based on summary. This procedure was applied to all selected instances across the chosen texts, using identical prompt templates for all languages. For non-English cases, we appended instructions specifying the target language of generation.
Controlling the generation process through explicit instructions is crucial for our study, as it ensures that the resulting LLM-generated persuasive texts remain semantically comparable to human-written texts. This comparability is essential for reliable evaluation of both persuasion detection performance and linguistic differences between human- and AI-generated persuasive texts.
All approaches for generating synthetic text with persuasion, along with example prompts, are included in Table 1 (details about our prompts presented in the Appendix B.3).
Approach Prompt Example
Paraphrased Persuasion Given a passage, please paraphrase it. The original content meaning and level of persuasion must be strictly preserved. The passage is: <passage>
Rewritten Persuasion (Subtle Effect) Given a passage, please rewrite it to make its persuasiveness more subtle and harder to detect. The original content and meaning should be the same. The passage is: <passage>
Rewritten Persuasion (Intensified Effect) Given a passage, please rewrite it so that the persuasion present in the content is refined for a stronger persuasive effect. The original content and meaning should be the same. The passage is: <passage>
Open-Ended Persuasion Given a sentence, please write a short piece of text. The short text must contain persuasion. The sentence is: <sentence>
Table 1: Overview of four approaches used for generating persuasive texts with LLMs. Each method represents a distinct level of control over persuasive strength and content nature.
3.2 Persuaficial Dataset Construction
Persuaficial is an AI-generated persuasion dataset constructed using multiple LLMs and diverse generation approaches. For Paraphrasing, Rewriting (Subtle Effect), and Rewriting (Intensified Effect), we sample 1,000 texts from three real-world persuasion datasets (described in Section 2). Each sample includes 500 texts annotated as persuasive and 500 labeled as non-persuasive [^3] The only exception was German: the corpus contained only 420 non-persuasive texts, so we included all of them and randomly sampled 580 persuasive texts.. Each selected persuasive text is treated as <passage> (see Table 1) and serves as input for the generation method. For Open-ended generation, we first summarize each selected persuasive text into a factual statements. We use the resulting <sentence> (see Table 1) for the generation of persuasive synthetic text.
We employ open-weight and proprietary LLMs for dataset construction. The open models include Gemma 3 27b it and Llama 3.3 70B. The commercial models are Gemini 2.0 Flash and GPT 4.1 Mini. Additional details on dataset creation and hyperparameter settings are provided in Appendix B. Details about the models, APIs used, and the rationale for model selection is available in Appendix C.
3.3 Persuaficial Quality Evaluation
Pre-Generation Quality Evaluation.
As mentioned, for the Open-ended generation setting, we first summarize each selected persuasive text into a short <sentence>. To ensure these <sentence>s accurately represent the source human text, we conducted a human evaluation following explicit and rigorous annotation guidelines (see Appendix D.1).
Two annotators were first trained by one of the authors, who has prior experience in annotation. A small training sample of 50 English sentences was selected, and the annotators independently applied the annotation guidelines, with opportunities to discuss their decisions. After completing the training phase, they reviewed and discussed their independent evaluations to align their understanding of the guidelines. The annotations from this training phase were excluded from further evaluation.
For the final evaluation, a sample of 200 English <sentence>s was selected. Two independent annotators assessed each <sentence> for factual correspondence. We then computed the accuracy of the LLM-generated <sentence>s, considering as positive only those instances where both annotators independently agreed that a sentence was factual. The resulting accuracy of the LLM generation process was about 91.2%, suggesting that LLMs may be effective at transforming texts into a short factual statements. Moreover, most mismatches between the generated <sentence>s and source texts were minor in nature, e.g., converting an exclamatory formulation (“Introduce the law!”) into a declarative one (“The law will be introduced.”). This result suggests that the generated factual sentences are of generally high quality and are unlikely to negatively impact the overall quality of the resulting Persuaficial dataset and our experiments.
Post-Generation Quality Evaluation.
To ensure that the final Persuaficial dataset meets the intended goal of containing persuasive content produced by LLMs under controlled prompting conditions, we conducted a multi-stage post-generation quality evaluation. While the pre-generation evaluation ensured that the factual sentences for Open-ended approach were valid, the post-generation evaluation verifies whether the LLM-generated persuasive variants are (1) faithful to the target factual content, (2) persuasive, and (3) faithful to the instruction from persuasion generation approach.
We adopted a two-layer rigorous verification design that separates basic validity checking from persuasion-specific judging. We verified 400 generated English texts, each independently annotated by two trained annotators following detailed instructions (Appendix D.2). As a result, we report an overall accuracy metric, defined as the proportion of generated texts unanimously annotated as valid by two annotators, where validity required that all three criteria received a positive annotation.
Due to the conservative requirement that all three criteria be jointly satisfied, the overall accuracy is 88.2%. Most invalid cases involve only minor factual deviations rather than substantive inconsistencies. When considering persuasion-related criteria alone, accuracy grows to 97.69%, indicating that LLMs reliably generate persuasive text and justifying the use of this data for subsequent comparisons between human- and AI-generated persuasive texts.
3.4 Data Statistics
For each language, we sampled 1,000 human-written passages from the original datasets, including 500 persuasive texts. For each persuasive example, we generated AI-counterparts using four LLMs and four generation approaches (4 models × 4 approaches = 16 generation configurations). This resulted in approx. 24,000 texts in English and 41,000 texts for the non-English languages. Overall, Persuaficial is a multilingual corpus of about 65,000 texts. Table 2 presents basic statistics for our dataset. Detailed statistics in Appendix E.
Data Statistic Human-written Paraphrasing Generation Rewriting Generation Open-ended Generation
Subtle Persuasion Intensive Persuasion
Average No. of Words 72 79 81 87 65
Average No. of Characters 452 538 568 605 450
Table 2: Basic statistics for human-written and for Persuaficial dataset, including LLM-generated persuasive texts.
4 Automatic Detection of Human and AI-Generated Persuasion
4.1 Experimental Setup
For our experiments, we use Persuaficial, which comprises artificially generated persuasive texts. Moreover, we use human-written counterparts. Each experiment uses data that is balanced across persuasive and non-persuasive classes.
For automatic persuasion detection, we employed four LLMs: GPT-4.1 Mini, Gemini 2.0 Flash, Gemma 3 27B Instruct, and Llama 3.3 70B. To ensure as deterministic outputs as possible, we set the temperature to 0 during classification. Since our goal is to detect persuasion, we formulate the task as a binary classification problem. All classifications were performed in a zero-shot setting. This approach aligns with our research question (RQ1). Moreover, studies show that zero-shot detection with modern LLMs (e.g., GPT-4) can outperform supervised models such as BERT on binary classification tasks Pelrine et al. (2023); Bang et al. (2023); Hassan and Lee (2020). Furthermore, Lucas et al. (2023) and Modzelewski et al. (2025) report that while fine-tuning BERT on multiple datasets results in poor generalization to unseen data, zero-shot LLMs maintain strong cross-domain performance. We evaluate persuasion detection performance using the F1 score. Further details supporting reproducibility, including the LLM classifiers setup and the prompt templates used for persuasion detection, are provided in Appendix F.
4.2 Results on English Datasets
Table 3 reports F1 scores for persuasion detection on three human-written balanced samples and their LLM-generated counterparts produced using four generation approaches.
On the Paraphrasing subset of our Persuaficial dataset, F1 scores are only marginally lower than those for human-written texts (on average 0.67% lower), indicating that paraphrasing preserves a similar level of difficulty for persuasion detection across human and generated texts. In contrast, Rewriting (Intensified) and Open-ended subsets yield the highest F1 scores. On average, persuasion is 9.75% easier to detect in open-ended scenario and 5.33% easier when persuasion is intensified. This makes open-ended generated persuasive texts the easiest setting for LLM-based detection. We hypothesize that models tend to over-express explicit persuasive cues when prompted to generate persuasive text freely or while intensifying persuasion, which in turn makes these texts more easily detectable. The opposite pattern emerges for Rewriting (Subtle persuasion), where F1 scores drop substantially, by 20.42% on average. This suggests that reducing overt persuasive markers makes persuasion significantly harder to detect, even for strong LLM detectors. Importantly, these patterns are highly consistent across datasets and across all detector models. This may indicate that the effects generalize across domains and are independent of the specific LLM used for detection.
In summary, addressing RQ1, the detectability of LLM-generated persuasive text depends on the generation approach: texts produced via open-ended and intensified persuasion are easier to detect, whereas subtly persuasive generations remain substantially more challenging for current LLM-based detectors. Detailed results in Appendix G.1.
Classifier Models Human-written Paraphrasing Generation Rewriting Generation Open-ended Generation
Subtle Persuasion Intensive Persuasion
Sample of Persuaficial generated based on: SemEval 2023 data
GPT 4.1 Mini 0.7398 $0.7007\downarrow$5% $0.4031\downarrow$46% $0.8148\uparrow$10% $0.8964\uparrow$21%
Llama 3.3 70B 0.7459 $0.7207\downarrow$3% $0.4577\downarrow$39% $0.8111\uparrow$9% $0.8741\uparrow$17%
[…]
Sample of Persuaficial generated based on: DIPROMATS 2024 data
GPT 4.1 Mini 0.7567 $0.7461\downarrow$1% $0.4962\downarrow$34% $0.8100\uparrow$7% $0.8666\uparrow$15%
Llama 3.3 70B 0.7471 $0.7362\downarrow$1% $0.5696\downarrow$24% $0.7860\uparrow$5% $0.8292\uparrow$11%
[…]
Gemini 2.0 Flash 0.7518 $0.7427\downarrow$1% $0.6664\downarrow$11% $0.7640\uparrow$2% $0.7680\uparrow$2%
Sample of Persuaficial generated based on: ChangeMyView data
GPT 4.1 Mini 0.6233 $0.6356\uparrow$2% $0.4906\downarrow$21% $0.6739\uparrow$8% $0.7148\uparrow$15%
[…]
Gemma 3 27b it 0.6644 $0.6708\uparrow$1% $0.6334\downarrow$5% $0.6809\uparrow$2% $0.6843\uparrow$3%
[…]
Table 3: F1 scores for persuasion detection on English data. The first column reports performance on human-annotated texts. The remaining columns show performance on LLM-generated texts. For generated data, each value represents the average F1 score obtained from classification of texts generated by four different LLMs. Detailed results without averaging F1 scores in Appendix G.1.
4.3 Results on Non-English Datasets
Table 4 shows persuasion detection results for German, French, Italian, Polish, and Russian. The patterns observed in English hold consistently across all languages and classifiers. Paraphrasing preserves a difficulty level similar to human-written texts, whereas intensified rewriting and open-ended generation yield the highest F1 scores. Open-ended generation frequently yields F1 scores above 0.9, indicating that persuasive text is easiest to detect in this setting. In contrast, subtle rewriting causes the largest drop in performance. These trends suggest that generation approaches influence persuasion detectability, with effects that may generalize across languages. Detailed results in Appendix G.2.
Classifier Models Human-written Paraphrasing Generation Rewriting Generation Open-ended Generation
Subtle Persuasion Intensive Persuasion
German
[…]
Llama 3.3 70B 0.7361 $0.7248\downarrow$2% $0.4398\downarrow$40% $0.8474\uparrow$15% $0.9345\uparrow$27%
Gemma 3 27b it 0.7655 $0.7763\uparrow$1% $0.6664\downarrow$13% $0.8512\uparrow$11% $0.9004\uparrow$18%
[…]
French
GPT 4.1 Mini 0.7505 $0.7454\downarrow$1% $0.4290\downarrow$43% $0.8456\uparrow$13% $0.9251\uparrow$23%
Llama 3.3 70B 0.7605 $0.7450\downarrow$2% $0.4527\downarrow$40% $0.8432\uparrow$11% $0.9172\uparrow$21%
[…]
Gemini 2.0 Flash 0.7812 $0.7860\uparrow$1% $0.6800\downarrow$13% $0.8314\uparrow$6% $0.8418\uparrow$8%
Italian
GPT 4.1 Mini 0.7471 $0.7330\downarrow$2% $0.4246\downarrow$43% $0.8428\uparrow$13% $0.9195\uparrow$23%
Llama 3.3 70B 0.7584 $0.7172\downarrow$5% $0.4285\downarrow$43% $0.8420\uparrow$11% $0.9161\uparrow$21%
Gemma 3 27b it 0.7659 $0.7804\uparrow$2% $0.6610\downarrow$14% $0.8399\uparrow$10% $0.8686\uparrow$13%
Gemini 2.0 Flash 0.7986 $0.7938\downarrow$1% $0.6781\downarrow$15% $0.8301\uparrow$4% $0.8408\uparrow$5%
Polish
GPT 4.1 Mini 0.7330 $0.7060\downarrow$4% $0.4580\downarrow$38% $0.8483\uparrow$16% $0.9367\uparrow$28%
Llama 3.3 70B 0.7676 $0.7389\downarrow$4% $0.5041\downarrow$34% $0.8518\uparrow$11% $0.9206\uparrow$20%
Gemma 3 27b it 0.7728 $0.7783\uparrow$1% $0.6919\downarrow$10% $0.8427\uparrow$9% $0.8834\uparrow$14%
[…]
Russian
GPT 4.1 Mini 0.7246 $0.7073\downarrow$2% $0.4392\downarrow$39% $0.8242\uparrow$14% $0.9017\uparrow$24%
Llama 3.3 70B 0.7408 $0.7164\downarrow$3% $0.4312\downarrow$42% $0.8324\uparrow$12% $0.9086\uparrow$23%
Gemma 3 27b it 0.7360 $0.7416\uparrow$1% $0.6128\downarrow$17% $0.8098\uparrow$10% $0.8562\uparrow$16%
Gemini 2.0 Flash 0.7683 $0.7616\downarrow$1% $0.6795\downarrow$12% $0.7877\uparrow$3% $0.8019\uparrow$4%
Table 4: F1 scores for persuasion detection on non-English data samples. The first column reports performance on human-annotated texts. The remaining columns show performance on LLM-generated texts. For generated data, each value represents the average F1 score obtained from classification of texts generated by four different LLMs. Detailed results without averaging F1 scores in Appendix G.2
5 Linguistic Differences Between Machine and Human Persuasion
In this section, we investigate the linguistic differences between human-written and AI-generated persuasive texts. We focus on English due to the limited availability of high-quality datasets in other languages. While the SemEval 2023 Task 3 data provides a multilingual resource Srba et al. (2024); Piskorski et al. (2023c), relying solely on it could introduce dataset-specific biases. To mitigate this, we use English part of Persuaficial and human-written counterparts from three well-established datasets.
5.1 Our Approach for Linguistic Analysis
To investigate the linguistic differences between human-written persuasive texts and LLM-generated persuasive texts, we adopt an explainable, feature-based analysis grounded in stylometry. Our objective is to identify the linguistic features that most strongly differentiate LLM-generated persuasive texts from human-written ones. For each linguistic feature, we compare the distributions of the two groups using effect-size-based analysis together with significance testing. Effect sizes quantify the magnitude of the difference between human and AI-generated texts for each feature Frey (2021), while significance testing evaluates whether these distributional differences are statistically meaningful. Our analysis identifies which linguistic properties differ systematically between human- and AI-produced persuasive texts.
5.2 Experimental Setup
We first represent each persuasive human text and its AI-generated counterpart using StyloMetrix Okulska et al. (2023). For each text, we calculate a 196-dimensional vector of linguistic features. This results in a tabular representation, where each row corresponds to a text encoded by its computed linguistic features. We utilize StyloMetrix, because it is an open-source tool that provides fully interpretable, linguistically grounded feature representations. Moreover, prior work has demonstrated its effectiveness for persuasion detection using classical machine learning models Modzelewski et al. (2023, 2024), confirming its suitability for analyzing the linguistic characteristics of persuasive texts. Finally, to further justify our choice, we show that StyloMetrix features with classical machine learning can distinguish human-written from AI-generated persuasive texts in Appendix H.
For each linguistic feature, we directly compare its distribution in human-written and LLM-generated persuasive texts, conducting this analysis separately for each generation approach. We utilize Cohen’s d statistic (see Appendix I for definiton and how we computed it in our experiments) which is a type of effect size measure used to represent the magnitude of differences between two groups on a given variable Frey (2021). To evaluate whether feature distributions significantly differ between human and AI-generated texts, we perform Wilcoxon signed-rank tests Wilcoxon (1945), a non-parametric paired test appropriate for comparing matched text pairs Peyrard et al. (2021); Dror et al. (2018). Wilcoxon signed-rank tests have seen widespread adoption in the NLP community Karwa and Singh (2025); Zhou et al. (2025); Ciaccio et al. (2025), including in studies comparing feature distributions Stodden and Kallmeyer (2020).
G_ACTIVE The proportion of verbs in the text used in the active voice. L_ADV_SUPERLATIVE Measures how often superlative adverbial (and some adjective-as-adverb) forms appear in a text. L_ADV_COMPARATIVE The proportion of tokens that are adverbs used in comparative degree (e.g., "more", "less", or marked comparative forms) L_FUNC_A The proportion of tokens in a text that are function words. L_CONT_T The proportion of unique content-word forms in relative to total tokens. L_CONT_A The proportion of tokens in a text that are content words. L_PUNCT_COM Comma incidence measures frequency of commas relative to text length. L_PUNCT_DASH Measures the density of dashes within a text. L_PUNCT_DOT Measures the incidence of periods (dots) relative to the total number of words. L_PLURAL_NOUNS Measures the density of plural nouns within a text. LTOKEN_RATIO_LEM (ST_TYPE_TOKEN_RATIO_LEMMAS) The ratio of unique lemmas to total tokens. POS_ADJ The proportion of tokens in the text that are adjectives, indicating the level of descriptiveness. POS_NOUN The proportion of tokens in the text that are nouns. POS_PRO The proportion of tokens in the text that are pronouns. PS_CAUSE Measures the incidence of linking words and phrases related to cause and purpose. SENT_D_NP (ST_SENT_D_NP) Measures the average proportion of noun phrase (NPs) tokens relative to sentence length, averaged over all sentences in a document. SENT_D_PP (ST_SENT_D_PP) Measures the average proportion of tokens that belong to prepositional phrases (PPs) in each sentence, averaged over all sentences in the document. SENT_D_VP (ST_SENT_D_VP) Measures the average proportion of tokens in a sentence that are not marked with a verb tense, relative to total sentence length, averaged over all sentences in the document. SENT_ST_DIFFER (ST_SENT_DIFFERENCE) Quantifies syntactic variation between consecutive sentences by comparing their dependency label sets and averaging over the document. SENT_ST_WPERSENT Indicates the normalized difference between the total number of tokens and the number of sentences in a document (a proxy for sentence length). ST_REPET_WORDS (ST_REPETITIONS_WORDS) Measures the level of lexical repetition by computing the proportion of repeated word tokens in a text normalized by total token count. SY_EXCLAMATION Measures the proportion of unique word tokens that appear in exclamatory sentences relative to all tokens in the text. SY_IMPERATIVE Measures the proportion of unique alphabetic words that appear in sentences classified as imperative, relative to all tokens in the document. SY_INV_PATTERNS (SY_INVERSE_PATTERNS) A syntactic feature that measures the frequency of inverted sentence structures within a text. SY_NARRATIVE Measures the proportion of tokens in declarative sentences relative to all tokens. VF_INFINITIVE A syntactic feature that measures the proportion of infinitive verb forms. VT_MIGHT Measures the frequency of "might" in a text.
Figure 1: StyloMetrix features with the highest discriminative importance in distinguishing human-written from LLM-generated persuasive text.
5.3 Results and Analysis
We computed Cohen’s d values for 196 linguistic features across four generation approaches and four LLMs utilized to generate texts for our Persuaficial dataset. Table 5 sorts top linguistic features by the absolute Cohen’s d ($|C_{d}|$) values per model and generation approach ($G_{P}$). Our analysis and discussion is based on the twenty features with the largest $|C_{d}|$ for each model and generation strategy (more detailed tables available in Appendix J). Statistical analysis using Wilcoxon signed-rank tests confirmed that all twenty features for each scenario exhibit significant distributional differences between human-written and AI-generated persuasive texts. Figure 1 provides definitions of the key differentiating features.
High values of features such as L_CONT_T (the proportion of unique content-word forms relative to total tokens), LTOKEN_RATIO_LEM (the ratio of unique lemmas to total tokens), and L_CONT_A (the proportion of tokens that are content words) indicate that AI-generated texts tend to contain more varied content words and higher informational density per sentence. These patterns suggest that lexical diversity and content richness are characteristic markers of AI authorship. Similarly, low values for ST_REPET_WORDS are indicative of AI-generated persuasive texts, suggesting that reduced word repetition serves as a strong signal of LLM-generated text. A higher proportion of function words (L_FUNC_A) indicates that a persuasive text is likely human-written. This means that frequent use of grammatical connectors (such as articles, prepositions, pronouns, and auxiliary verbs) is a signal of human text. Furthermore, AI-generated persuasive texts generally exhibit lower punctuation density, especially in texts from Llama and GPT-4.1-mini. However, certain punctuation marks, including commas (L_PUNCT_COM) and dashes (L_PUNCT_DASH), occur more frequently in these AI texts. The rarity of syntactically marked constructions, such as inversions (SY_INV_PATTERNS), is a distinguishing feature of AI text, as these complex syntactic patterns are more typical of human-written persuasive texts.
In AI-generated texts that intensify persuasion, comparative and superlative adverbs (L_ADV_COMPARATIVE with words like "more", "faster", and L_ADV_SUPERLATIVE with words like "best", "worst") appear more frequently than in human-written texts. This suggests that AI strengthens persuasive language through the increased use of adverbial modifiers, highlighting a distinctive stylistic strategy in LLM-generated intensified texts.
In AI-generated texts that aim to make persuasion more subtle, modal verbs such as "might" (VT_MIGHT) occur more frequently than in human-written texts. Similarly, narrative framing (SY_NARRATIVE), defined as the proportion of tokens in declarative sentences relative to all tokens, is more prevalent in AI subtle rewritings. These patterns may indicate that AI softens persuasion by using modal hedges and favors neutral declarative constructions over exclamatory sentences or rhetorical questions more often than humans do.
Open-ended AI-generated texts exhibit a consistent linguistic profile characterized by high lexical diversity and elevated content-word density (e.g., L_CONT_T, L_CONT_A, LTOKEN_RATIO_LEM). AI systems also show substantially lower function-word usage and reduced lexical repetition. In addition, AI-generated texts rely more heavily on imperative and infinitival constructions while avoiding marked syntactic patterns such as inversions, which may result in structurally simpler and more schematic syntax.
Paraphrasing Rewriting for Subtle Effects Rewriting for Intensified Effects Open-ended
Feature Name $C_{d}$ Feature Name $C_{d}$ Feature Name $C_{d}$ Feature Name $C_{d}$
Generating Model: GPT 4.1 Mini
L_CONT_T 0.51 L_CONT_T 0.70 L_CONT_T 0.77 L_CONT_T 1.62
L_CONT_A 0.46 L_CONT_A 0.62 L_CONT_A 0.77 L_CONT_A 1.43
SY_INV_PATTERNS -0.45 VT_MIGHT 0.62 L_PUNCT_DASH 0.72 LTOKEN_RATIO_LEM 1.32
LTOKEN_RATIO_LEM 0.37 SY_INV_PATTERNS -0.61 L_FUNC_A -0.57 SENT_D_NP 1.05
L_FUNC_A -0.36 L_PLURAL_NOUNS 0.54 LTOKEN_RATIO_LEM 0.51 ST_REPET_WORDS -1.00
Generating Model: Llama 3.3 70B
SENT_ST_WPERSENT 0.71 SY_INV_PATTERNS -0.75 SENT_ST_WPERSENT 0.80 VF_INFINITIVE 1.34
SENT_ST_DIFFER -0.70 G_ACTIVE -0.72 L_ADJ_POSITIVE 0.72 SY_IMPERATIVE 1.24
L_PUNCT_COM 0.67 SENT_ST_WPERSENT 0.70 L_PUNCT_COM 0.71 G_ACTIVE -0.91
SY_INV_PATTERNS -0.61 L_CONT_T 0.62 L_CONT_T 0.66 SENT_D_VP 0.83
L_CONT_T 0.54 SENT_D_PP 0.62 POS_ADJ 0.65 L_PUNCT_DOT -0.81
Generating Model: Gemma 3 27b it
L_CONT_T 0.75 L_CONT_T 1.02 L_CONT_T 1.04 SY_IMPERATIVE 1.74
L_CONT_A 0.67 L_CONT_A 1.02 L_CONT_A 1.01 L_CONT_T 1.4
L_FUNC_A -0.62 L_FUNC_A -0.89 L_FUNC_A -0.97 L_CONT_A 1.25
SY_INV_PATTERNS -0.61 POS_PRO -0.73 L_ADJ_POSITIVE 0.90 SENT_D_NP 1.16
L_ADV_SUPERLATIVE 0.43 SY_INV_PATTERNS -0.71 POS_ADJ 0.83 PS_CAUSE -1.15
Generating Model: Gemini 2.0 Flash
L_CONT_A 0.73 L_CONT_A 0.92 L_CONT_A 1.02 L_CONT_T 1.57
L_CONT_T 0.71 L_CONT_T 0.92 L_CONT_T 0.99 L_CONT_A 1.37
SY_INV_PATTERNS -0.58 L_FUNC_A -0.73 L_FUNC_A -0.82 SY_IMPERATIVE 1.32
L_FUNC_A -0.54 SY_INV_PATTERNS -0.72 L_ADJ_POSITIVE 0.78 LTOKEN_RATIO_LEM 1.20
L_PUNCT_COM 0.49 POS_NOUN 0.59 POS_ADJ 0.72 SY_EXCLAMATION 1.11
Table 5: Top linguistic features by absolute Cohen’s d from four generation approaches for all generating LLMs.
6 Related Work
Research on the persuasive capabilities of generative AI spans multiple disciplines, including computer science as well as the social and complexity sciences Duerr and Gloor (2021). Recent progress in large language models has drawn attention to their potential for persuasion and related applications Jin et al. (2024); Rogiers et al. (2024). Early studies by Wang et al. (2019) explored personalized persuasive dialogue systems designed to promote socially beneficial outcomes. Subsequent studies have investigated how individuals respond to persuasive machine-generated text and how they perceive its effectiveness Karinshak et al. (2023); Bai et al. (2025); Goldstein et al. (2024). In addition, Schoenegger et al. (2025) explored whether LLMs can be more persuasive than humans, while Pauli et al. (2025) analyzed the extent to which LLMs are capable of generating persuasive language.
Parallel work has investigated approaches for detecting persuasion. Da San Martino et al. (2019) presented corpus of news annotated at the fragment level with 18 persuasive techniques and proposed multi-granularity neural network to detect persuasion. Piskorski et al. (2023d) extended the taxonomy of techniques proposed by Da San Martino et al. (2019) and presented a multilingual dataset. Moreover, persuasion detection was a task of different recognized workshops such SemEval or SlavicNLP Da San Martino et al. (2020); Dimitrov et al. (2021); Piskorski et al. (2023b, 2025).
To the best of our knowledge, no prior work has investigated whether LLM-generated persuasive texts are easier to automatically detect than human-written ones. Furthermore, existing research has not provided a linguistic analysis comparing LLM-generated and human persuasive content.
7 Conclusion
In this work, we introduce Persuaficial, a multilingual benchmark of LLM-generated persuasive texts, comprising about 65,000 instances across English, German, French, Italian, Polish, and Russian. Our experiments show that the detectability of persuasion in generated texts strongly depends on the generation approaches: open-ended and rewriting with intensified persuasion increase detectable cues, whereas rewriting with subtle persuasion substantially reduces detection performance. These trends are consistent across languages and classifier models, indicating that generation prompts may shape the difficulty of persuasive text detection.
Through a detailed analysis, we identify key linguistic differences between human and AI-generated persuasive texts. AI-generated texts tend to exhibit higher lexical diversity, increased content-word density, and lower function-word usage, while complex syntactic patterns are more characteristic of human persuasive writing. Text generation approaches further modulate these features: intensified persuasion amplifies adverbial modifiers, whereas subtle persuasion relies on modal hedges and declarative constructions.
Overall, our findings demonstrate that LLM-generated persuasive texts are not only linguistically distinct from human-written texts but also vary in detectability depending on the generation approach. Persuaficial provides a valuable resource for future research on automated persuasion detection, cross-lingual NLP, and the study of linguistic differences between human and AI-generated persuasive content.
8 Limitations
While Persuaficial offers a large and diverse resource for studying AI persuasive text detection, several limitations remain. First, although the corpus covers six languages, our linguistic analysis focuses only on English. This decision stems from the limited availability of high-quality, human-written persuasive datasets in other languages and from the need for a controlled, comparable setup across human and AI-generated texts. Conducting the analysis exclusively on English avoids introducing dataset-specific biases that would arise from relying on a single non-English persuasion dataset, ensuring that the linguistic findings are not driven by characteristics of a particular corpus.
In our analysis, all classifier evaluations are conducted in a zero-shot setting, which aligns with the goals and research question of this work. We chose not to explore few-shot or alternatives, leaving these as directions for future research. Moreover, prior studies show that modern LLMs in zero-shot mode (e.g., GPT-4) can outperform supervised models such as BERT on binary classification tasks Pelrine et al. (2023); Bang et al. (2023); Hassan and Lee (2020), and that fine-tuned BERT models may generalize poorly to out-of-domain data compared to zero-shot LLMs Lucas et al. (2023); Modzelewski et al. (2025).
9 Ethics
Dataset
The Persuaficial dataset consists of synthetic persuasive texts generated by large language models for research purposes. To construct the dataset, we relied exclusively on three established human-authored persuasion datasets that are either publicly available or were used with explicit permission from their original authors. No personally identifiable information is included in the generated persuasive content, and no attempt is made to identify or infer the authorship of individual texts.
All human-written and synthetic texts were used solely for academic research on persuasion detection and linguistic analysis. The generation process preserves the semantic content of the source material while producing novel text, thereby avoiding the reproduction of identifiable original passages. To promote transparency, reproducibility, and responsible reuse, the Persuaficial dataset will be released under the Creative Commons Attribution 4.0 (CC BY 4.0) license.
Crowdsourcing was not used at any stage of dataset creation or validation. All individuals involved in verifying the quality of Persuaficial were researchers or trained annotators with prior experience in persuasion or manipulation annotation. The verification process was conducted independently and remained free from political, institutional, or commercial influence.
Computational resources
The use of large language models is associated with non-trivial computational and environmental costs Strubell et al. (2019). In this work, we mitigate these costs by avoiding model training or fine-tuning and relying exclusively on inference with pre-existing models. All experiments were conducted through third-party APIs, and we did not directly manage or allocate the underlying computational infrastructure. As a result, the overall computational footprint of this study was limited to inference-time usage.
Appendix A Human Dataset - Rationale Behind Our Choice
In our experiments, we adopt the concise definition of persuasion proposed by Piskorski et al. (2023d, a): “Persuasive text is characterized by a specific use of language in order to influence the reader”. We rely on this definition because it underpins the annotation guidelines of the SemEval 2023 Task 3 dataset, the largest publicly available resource for studying persuasion and one of the three human-written datasets used in our study. Consequently, we selected additional datasets that align well with this conceptualization of persuasion.
DIPROMATS 2024 (Moral et al., 2024), which build on the SemEval task, offers data that is directly compatible with this definition and is therefore suitable for our experiments. Finally, Piskorski et al. (2025) demonstrate that this definition can be effectively and reliably applied to debates and discussions, motivating our choice of the ChangeMyView dataset, comprising message exchanges between persuaders and persuadees, as an additional human-written source.
Appendix B Persuaficial Generation Process
B.1 Human-written Text Sample Selection
For each of the eight source datasets (DIPROMATS 2024, ChangeMyView, and SemEval 2023 Task 3 with six languages), we generated samples from a base of 500 persuasive human-written texts. By applying four generation approaches across four different LLMs, we produced 16 distinct machine-generated counterparts for every source text. Moreover, each sample contains 500 non-persuasive human-written texts.
B.2 Experimental Setup for LLM Persuasion Generation Process
We employed four Large Language Models. The open models include Gemma 3 27b it and Llama 3.3 70B. The commercial models are Gemini 2.0 Flash and GPT 4.1 Mini. To encourage more diverse, creative and less repetitive phrasing in the model outputs, we set the generation temperature to 0.8. Our choice of temperature for generating synthetic persuasive texts was directly informed by the settings used by Chen and Shu (2023) in a related task involving misinformation generation.
B.3 Prompt Templates for Persuaficial Dataset Creation
Figures 2, 3, 4, and 5 show prompt templates used during the LLM persuasion generation process. In our prompts, we adopt the concise definition of persuasion proposed by Piskorski et al. (2023d, a): “Persuasive text is characterized by a specific use of language in order to influence the reader”.
Generation approach: Paraphrasing Generation prompt System Prompt: You are an assistant helping researchers study persuasion. Your task is to paraphrase provided texts so that the meaning, and original persuasion techniques are strictly preserved. Do not make the provided texts less or more persuasive. Keep the same level of persuasion. Definition of persuasion is as follows: Persuasive text is characterized by a specific use of language in order to influence readers. We distinguish the following high-level persuasion approaches: Attack on reputation, Justification, Simplification, Distraction, Call, Manipulative wording. You will receive the original passage. Then, paraphrase the passage, ensuring the paraphrased version contains the exact same level of persuasion and meaning. Your final output should only be the paraphrased text. User Prompt: Only paraphrase the following passage in <language>
The passage is: <Human-written, persuasive text to paraphrase>
Keep the same level of persuasion. Provide the paraphrased text in <language>, in valid JSON format: { "generated_text":"Your paraphrased text in <language> here." }
Figure 2: Prompt template used for persuasive texts generation with LLMs using the Paraphrasing Generation approach.
Generation approach: Rewriting Generation - Subtle Persuasion prompt System Prompt: You are an assistant helping researchers study persuasion. Your task is to rewrite provided texts so that the persuasion is more subtle and harder to detect, while strictly preserving all original persuasion techniques. Definition of persuasion is as follows: Persuasive text is characterized by a specific use of language in order to influence readers. We distinguish the following high-level persuasion approaches: Attack on reputation, Justification, Simplification, Distraction, Call, Manipulative wording. You will receive the original passage. Then, rewrite the passage, ensuring the rewritten version contains the exact meaning. Your final output should only be the rewritten text. User Prompt: Only rewrite the following passage in <language> so that the persuasion is more subtle and harder to detect.
The passage is: <Human-written, persuasive text to rewrite>
Provide the rewritten text in <language>, in valid JSON format: { "generated_text":"Your rewritten text in <language> here." }
Figure 3: Prompt template used for persuasive texts generation with LLMs using the Rewriting Generation - Subtle Persuasion approach.
Generation approach: Rewriting Generation - Intensive Persuasion prompt System Prompt: You are an assistant helping researchers study persuasion. Your task is to rewrite provided texts so that the persuasion is refined for stronger persuasive effect, while strictly preserving all original persuasion techniques. Definition of persuasion is as follows: Persuasive text is characterized by a specific use of language in order to influence readers. We distinguish the following high-level persuasion approaches: Attack on reputation, Justification, Simplification, Distraction, Call, Manipulative wording. You will receive the original passage. Then, rewrite the passage, ensuring the rewritten version contains the exact meaning. Your final output should only be the rewritten text. User Prompt: Only rewrite the following passage in <language> so that the persuasion is refined for stronger persuasive effect.
The passage is: <Human-written, persuasive text to paraphrase>
Provide the rewritten text in <language>, in valid JSON format: { "generated_text":"Your rewritten text in <language> here." }
Figure 4: Prompt template used for persuasive texts generation with LLMs using the Rewriting Generation - Intensive Persuasion approach.
Generation approach: Open-ended Generation prompt System Prompt: You are an assistant helping researchers study persuasion. Your task is to generate a short text based on a provided passage. The short text must contain persuasion. Definition of persuasion is as follows: Persuasive text is characterized by a specific use of language in order to influence readers. We distinguish the following high-level persuasion approaches: Attack on reputation, Justification, Simplification, Distraction, Call, Manipulative wording. Your final output should only be the generated text. User Prompt: Generate a text in <language> based on the following passage in <language>. The passage is: <Summarized, factual, and non-persuasive input sentence> The generated text must contain persuasion. Provide the generated text in valid JSON format: { "generated_text":"Your generated text in <language> here." }
Obtaining a summarized  factual input sentence prompt System Prompt: You are a journalist assistant. Your task is to convert the provided text passage into a direct, single-sentence text. Do not add context such as ’The speaker said…’, ’The passage is about…’, ’The statement suggests…’, etc. Keep the meaning intact but make it stand alone. Do not add any additional information or actors. User Prompt: Restate the following passage in <language> as a single-sentence, neutral text in <language>. The passage is: <Human-written, persuasive text to summarize> Return in valid JSON format: { "generated_text":"Your restated sentence in <language> here." }
Figure 5: Prompt template used for persuasive texts generation with LLMs using the Open-ended Generation approach along with the prompt template used to obtain a summarized, factual, and non-persuasive sentence input from human-written persuasive texts.
Binary Detection of Persuasion prompt System Prompt: You are an assistant who detects persuasion in text. Persuasive text is characterized by a specific use of language in order to influence readers. We distinguish the following high-level persuasion approaches: Attack on reputation, Justification, Simplification, Distraction, Call, Manipulative wording. You are the expert who detects high-level persuasion. User Prompt: Analyze the following passage: <Text to analyze> Decide if the passage contains persuasion. You are very conservative in your final decisions and when you are not fully sure you answer ’No’. Do not provide any additional text, just JSON. Give only your final answer ’Yes’ or ’No’ in valid JSON format: { "decision": "’Yes’ if passage contains persuasion, ’No’ otherwise." }
Figure 6: Prompt template used for binary classification of persuasive texts with LLMs.
Appendix C Details on LLMs used in Experiments and selection rationale.
In our experiments, we employed four state-of-the-art LLMs: GPT-4.1 Mini, Gemini 2.0 Flash, Gemma 3 27B-IT, and Llama 3.3 70B. Our selection aimed to cover widely recognized, high-performing models while balancing accessibility and cost. Additionally, we included two open-weight models to provide experiments that can be reproduced without reliance on closed API-based models. Table 6 summarizes the LLMs used, including their knowledge cutoff dates, access methods, licenses, and model sizes.
API Model Name Knowledge Cutoff Date Access Details License Model Size
gemini-2.0-flash June 2024 Google API 07.2025 Commercial Not Disclosed
gpt-4.1-mini-2025-04-14 June 2024 OpenAI API 07.2025 Commercial Not Disclosed
meta-llama/Llama-3.3-70B-Instruct December 2023 DeepInfra API 07.2025 Meta Llama 3 Community 70B
google/gemma-3-27b-it August 2024 DeepInfra API 07.2025 Gemma Terms of Use 27B
Table 6: Large Language Models used in our experiments.
Appendix D Annotation Guidelines for Dataset Evaluation
D.1 Annotation Guidelines for Sentences Verification
Purpose of the Annotation Task.
The goal of this annotation task is to evaluate whether each LLM-generated <sentence> accurately reflects content present in its corresponding source human text. Annotators must independently judge whether the <sentence> faithfully reflects information explicitly stated in the source text, without adding, or altering factual content.
General Annotation Procedure.
  1. 1.
    Read the source persuasive human text in full to understand its factual content and context.
  2. 2.
    Read the generated <sentence> carefully and evaluate it against the factual correspondence.
  3. 3.
    Assign one binary label:
    Factual? Yes (1) / No (0)
  4. 4.
    Do not consider any stylistic preferences, or grammar.
Annotators should make decisions independently, without discussing individual cases during the evaluation phase.
Factual Correspondence Annotation.
  1. 1.
    All information in the <sentence> is explicitly stated in the source text.
    • No invented facts.
    • The <sentence> does not introduce generalizations (e.g., Fact present in a source text: "Adam Smith fainted after COVID-19 vaccination" $\rightarrow$ invalid <sentence>: "People fainted after COVID-19 vaccination")
    • No added assumptions or interpretations.
  2. 2.
    No main factual information from the source text is omitted in a way that distorts meaning.
  3. 3.
    The <sentence> is neutral and descriptive - Its purpose must be to summarize factual content, not to evaluate, interpret, or advise.
  4. 4.
    Statements must be verifiable based solely on the source text. Annotators should not use outside knowledge.
Examples of factual errors (should be labeled "No"):
  • Adding additional events or statistics not in the source
  • Reframing a claim as a fact (e.g., converting someone’s opinion into an asserted truth)
  • Omitting a main fact presented in source text that changes meaning.
D.2 Annotation Guidelines for Persuaficial Dataset Evaluation
These guidelines describe the annotation protocol for evaluating LLM-generated persuasive texts in the Persuaficial dataset. Each generated text is independently annotated by two annotators. The post-generation evaluation focuses on three key dimensions:
  • Factual Correspondence: Is the generated text faithful to the target factual content?
  • Persuasiveness: Does the text contain genuine persuasion?
  • Instruction Adherence: Does the text follow the specific persuasion instruction for its generation approach?
Factual Correspondence.
Goal of this step is to ensure the generated text accurately reflects the source content.
Instructions:
  • Open-Ended Generation: Refer to the factual sentence.
  • Paraphrasing / Rewriting Approaches: Refer to the original passage.
Assessment:
  • Valid (represented as 1): Text preserves the factual meaning of the source without introducing errors or contradictions.
  • Invalid (represented as 0): Text contains factual inaccuracies, omissions, or misrepresentations.
Note: Only factual distortion triggers an Invalid label.
Persuasiveness
The generated text must contain any persuasive elements.
For this task, we define persuasive text as text characterized by a specific use of language in order to influence readers Piskorski et al. (2023a, d). The generated text must be labeled as persuasive (represented as 1) if it exhibits any of the following high-level persuasion strategies:
  • Attack on reputation: the argument does not address the topic itself, but targets the participant (personality, experience, deeds, etc.) in order to question and/or to undermine his credibility. The object of the argumentation can also refer to a group of individuals, an organization, an object, or an activity,
  • Justification: the argument is made of two parts, a statement and an explanation or appeal, where the latter is used to justify and/or to support the statement,
  • Simplification: the argument excessively simplifies a problem, usually regarding the cause, the consequence, or the existence of choices,
  • Distraction: the argument takes focus away from the main topic or argument to distract the reader,
  • Call: the text is not an argument but an encouragement to act or to think in a particular way,
  • Manipulative wording: the text is not an argument per se, but uses specific language, which contains words or phrases that are either non-neutral, confusing, exaggerating, loaded, etc., in order to impact the reader emotionally.
If any of these strategies are present, the sentence must be labeled 1 (persuasive) for the persuasiveness criterion.
Instruction Adherence.
The goal is to verify that the text aligns with the intended generation approach.
Instructions for Annotators:
  1. 1.
    Compare the generated text to the prompt provided to the model.
  2. 2.
    Label Compliant (represented as 1) if the text follows the prompt goal; Non-Compliant (represented as 0) if it deviates.
Appendix E Persuaficial Dataset - Additional Statistics
Table 7 summarizes the basic statistics of both human-written and LLM-generated texts in the Persuaficial dataset. The table reports average word, average characters and number of words across all languages and generation types. Moreover, we show statistics in general for full Persuaficial dataset.
$Avg_{w}$ $Avg_{ch}$ Count
dataset type
Persuaficial Rewriting (intensified) 87 605 16320
Open-ended 65 450 16320
Paraphrasing 81 538 16320
Rewriting (subtle) 87 568 16320
English Human 117 695 3000
Rewriting (intensified) 118 791 6000
Open-ended 60 391 6000
Paraphrasing 111 727 6000
Rewriting (subtle) 110 742 6000
French Human 46 288 1000
Rewriting (intensified) 75 498 2000
Open-ended 76 498 2000
Paraphrasing 66 430 2000
Rewriting (subtle) 69 465 2000
German Human 44 314 1000
Rewriting (intensified) 68 512 2320
Open-ended 71 524 2320
Paraphrasing 59 434 2320
Rewriting (subtle) 63 476 2320
Italian Human 48 313 1000
Rewriting (intensified) 75 511 2000
Open-ended 75 507 2000
Paraphrasing 66 443 2000
Rewriting (subtle) 70 488 2000
Polish Human 46 327 1000
Rewriting (intensified) 70 526 2000
Open-ended 61 454 2000
Paraphrasing 62 459 2000
Rewriting (subtle) 66 493 2000
Russian Human 40 285 1000
Rewriting (intensified) 55 417 2000
Open-ended 58 442 2000
Paraphrasing 48 360 2000
Rewriting (subtle) 52 400 2000
Table 7: Basic statistics for human-written and for LLM-generated persuasive texts in our Persuaficial dataset. We present basic statistics for general full dataset, but also on samples that present all languages. $Avg_{w}$ stands for average words and $Avg_{ch}$ stands for average characters.
Appendix F Persuasion Detection - Experimental Setup Details
F.1 Evaluation Text Sample Creation
Our evaluation framework comprised a total of 68 distinct classification experiments for each of the eight source datasets (eight as SemEval data can be counted as 6 datasets each in different language). This setup involved testing every combination of four generation approaches and four generating LLMs. The resulting 16 sets of machine-generated text, along with a baseline of human-written persuasive text, were then evaluated by four different classifying LLMs, leading to the (16 + 1) × 4 = 68 experimental conditions.
Our evaluation framework is designed to isolate the impact of these generated texts. Each experiment’s evaluation set is composed of two halves:
  • A constant set of 500 human-written, non-persuasive texts from the original dataset. Predictions for this set were calculated once for each model and reused across all experiments for that dataset.
  • A variable set of 500 persuasive texts, which consists of the LLM-generated samples for a given generation approach.
Consequently, any variation in the F1 score between different generation models on the same dataset is attributable solely to the model’s performance on the generated persuasive samples. For our baseline experiments, labeled ’Human-written’, these persuasive samples are the original human-written texts from the dataset.
F.2 Experimental Setup for LLM Classification Process
We employed four Large Language Models. The open models include Gemma 3 27b it and Llama 3.3 70B. The commercial models are Gemini 2.0 Flash and GPT 4.1 Mini. To ensure determinism in the classification predictions, we set the classification temperature to 0.0. Rationale for models selection provided in Appendix C.
F.3 Prompt Templates for Persuasion Detection
Figure 6 shows a prompt template used during the LLM persuasion detection process. In our prompts, we adopt the concise definition of persuasion proposed by Piskorski et al. (2023d, a): “Persuasive text is characterized by a specific use of language in order to influence the reader”.
Appendix G Human vs. AI-generated Persuasion Detection - Detailed Results
G.1 Detailed results on English texts of Persuaficial dataset
In this appendix, we present detailed F1 scores for persuasion detection across different subsets of the Persuaficial datasets and their human-written counterparts. Results are presented for each LLM generation strategy. Table 8 reports results for SemEval 2023 Task 3 texts and their AI-generated counterparts, Table 9 for the DIPROMATS 2024 dataset, and Table 10 for the ChangeMyView dataset. For each dataset, classifier performance on human-written texts (first column) is compared with performance on LLM-generated texts produced via paraphrasing, rewriting with subtle or intensive persuasion, and open-ended generation. Results are further broken down by both the generating model and the classifier model, highlighting how different generation approaches influence the detectability of persuasion.
Classifier Models Human-written Paraphrasing Generation Rewriting Generation Open-ended Generation
Subtle Persuasion Intensive Persuasion
Generating model: GPT 4.1 Mini
GPT 4.1 Mini 0.7398 $0.6984\downarrow$6% $0.3837\downarrow$48% $0.7638\uparrow$3% $0.8969\uparrow$21%
Llama 3.3 70B 0.7459 $0.7310\downarrow$2% $0.4444\downarrow$40% $0.7757\uparrow$4% $0.8741\uparrow$17%
[…]
Gemini 2.0 Flash 0.7551 $0.7487\downarrow$1% $0.6407\downarrow$15% $0.7780\uparrow$3% $0.8117\uparrow$7%
Generating model: Llama 3.3 70B
GPT 4.1 Mini 0.7398 $0.6831\downarrow$8% $0.4411\downarrow$40% $0.7870\uparrow$6% $0.8969\uparrow$21%
Llama 3.3 70B 0.7459 $0.6911\downarrow$7% $0.4811\downarrow$36% $0.7791\uparrow$4% $0.8741\uparrow$17%
Gemma 3 27b it 0.7572 $0.7479\downarrow$1% $0.6746\downarrow$11% $0.8082\uparrow$7% $0.8562\uparrow$13%
Gemini 2.0 Flash 0.7551 $0.7476\downarrow$1% $0.6691\downarrow$11% $0.7921\uparrow$5% $0.8117\uparrow$7%
Generating model: Gemma 3 27b it
GPT 4.1 Mini 0.7398 $0.7025\downarrow$5% $0.3445\downarrow$53% $0.8611\uparrow$16% $0.8969\uparrow$21%
Llama 3.3 70B 0.7459 $0.7222\downarrow$3% $0.4118\downarrow$45% $0.8469\uparrow$14% $0.8741\uparrow$17%
Gemma 3 27b it 0.7572 $0.7675\uparrow$1% $0.6241\downarrow$18% $0.8383\uparrow$11% $0.8562\uparrow$13%
Gemini 2.0 Flash 0.7551 $0.7614\uparrow$1% $0.6202\downarrow$18% $0.8039\uparrow$6% $0.8117\uparrow$7%
Generating model: Gemini 2.0 Flash
GPT 4.1 Mini 0.7398 $0.7188\downarrow$3% $0.4430\downarrow$40% $0.8472\uparrow$15% $0.8949\uparrow$21%
Llama 3.3 70B 0.7459 $0.7385\downarrow$1% $0.4936\downarrow$34% $0.8428\uparrow$13% $0.8741\uparrow$17%
Gemma 3 27b it 0.7572 $0.7652\uparrow$1% $0.6613\downarrow$13% $0.8362\uparrow$10% $0.8562\uparrow$13%
[…]
Table 8: F1 scores for persuasion detection on English data sample of Persuaficial. More specifically, on sample of Persuaficial generated using English texts from SemEval 2023 Task 3 dataset. The first column reports performance on English texts from SemEval 2023 Task 3 human-annotated texts. The remaining columns show performance on LLM-generated English counterparts.
Classifier Models Human-written Paraphrasing Generation Rewriting Generation Open-ended Generation
Subtle Persuasion Intensive Persuasion
Generating model: GPT 4.1 Mini
GPT 4.1 Mini 0.7567 $0.7435\downarrow$2% $0.4948\downarrow$35% $0.7866\uparrow$4% $0.8666\uparrow$15%
Llama 3.3 70B 0.7471 $0.7348\downarrow$2% $0.5795\downarrow$22% $0.7679\uparrow$3% $0.8292\uparrow$11%
[…]
Gemini 2.0 Flash 0.7518 $0.7449\downarrow$1% $0.6711\downarrow$11% $0.7595\uparrow$1% $0.7680\uparrow$2%
Generating model: Llama 3.3 70B
GPT 4.1 Mini 0.7567 $0.7338\downarrow$3% $0.5595\downarrow$26% $0.7967\uparrow$5% $0.8666\uparrow$15%
Llama 3.3 70B 0.7471 $0.7314\downarrow$2% $0.6251\downarrow$16% $0.7775\uparrow$4% $0.8292\uparrow$11%
Gemma 3 27b it 0.7473 $0.7410\downarrow$1% $0.6928\downarrow$7% $0.7749\uparrow$4% $0.7994\uparrow$7%
Gemini 2.0 Flash 0.7518 $0.7400\downarrow$2% $0.7002\downarrow$7% $0.7652\uparrow$2% $0.7680\uparrow$2%
Generating model: Gemma 3 27b it
GPT 4.1 Mini 0.7567 $0.7672\uparrow$1% $0.3951\downarrow$48% $0.8404\uparrow$11% $0.8666\uparrow$15%
[…]
Gemini 2.0 Flash 0.7518 $0.7459\downarrow$1% $0.6232\downarrow$17% $0.7662\uparrow$2% $0.7680\uparrow$2%
Generating model: Gemini 2.0 Flash
GPT 4.1 Mini 0.7567 $0.7399\downarrow$2% $0.5353\downarrow$29% $0.8163\uparrow$8% $0.8666\uparrow$15%
Llama 3.3 70B 0.7471 $0.7336\downarrow$2% $0.5838\downarrow$22% $0.7911\uparrow$6% $0.8292\uparrow$11%
[…]
Gemini 2.0 Flash 0.7518 $0.7400\downarrow$2% $0.6711\downarrow$11% $0.7652\uparrow$2% $0.7680\uparrow$2%
Table 9: F1 scores for persuasion detection on English data sample of Persuaficial. More specifically, on sample of Persuaficial generated using DIPROMATS 2024 dataset. The first column reports performance on DIPROMATS 2024 human-annotated texts. The remaining columns show performance on LLM-generated counterparts.
Classifier Models Human-written Paraphrasing Generation Rewriting Generation Open-ended Generation
Subtle Persuasion Intensive Persuasion
Generating model: GPT 4.1 Mini
GPT 4.1 Mini 0.6233 $0.6337\uparrow$2% $0.4941\downarrow$21% $0.6582\uparrow$6% $0.7148\uparrow$15%
[…]
Gemma 3 27b it 0.6644 $0.6745\uparrow$2% $0.6388\downarrow$4% $0.6809\uparrow$2% $0.6836\uparrow$3%
[…]
Generating model: Llama 3.3 70B
GPT 4.1 Mini 0.6233 $0.6137\downarrow$2% $0.4619\downarrow$26% $0.6481\uparrow$4% $0.7148\uparrow$15%
Llama 3.3 70B 0.6517 $0.6307\downarrow$3% $0.5226\downarrow$20% $0.6564\uparrow$1% $0.6831\uparrow$5%
[…]
Gemini 2.0 Flash 0.6671 $0.6607\downarrow$1% $0.6112\downarrow$8% $0.6717\uparrow$1% $0.6770\uparrow$1%
Generating model: Gemma 3 27b it
GPT 4.1 Mini 0.6233 $0.6572\uparrow$5% $0.4840\downarrow$22% $0.7009\uparrow$12% $0.7148\uparrow$15%
Llama 3.3 70B 0.6517 $0.6574\uparrow$1% $0.5515\downarrow$15% $0.6758\uparrow$4% $0.6831\uparrow$5%
Gemma 3 27b it 0.6644 $0.6754\uparrow$2% $0.6349\downarrow$4% $0.6836\uparrow$3% $0.6845\uparrow$3%
[…]
Generating model: Gemini 2.0 Flash
GPT 4.1 Mini 0.6233 $0.6379\uparrow$2% $0.5226\downarrow$16% $0.6885\uparrow$10% $0.7148\uparrow$15%
[…]
Gemma 3 27b it 0.6644 $0.6690\uparrow$1% $0.6465\downarrow$3% $0.6818\uparrow$3% $0.6845\uparrow$3%
[…]
Table 10: F1 scores for persuasion detection on English data sample of Persuaficial. More specifically, on sample of Persuaficial generated using ChangeMyView dataset. The first column reports performance on ChangeMyView human-annotated texts. The remaining columns show performance on LLM-generated counterparts.
G.2 Detailed results on non-English texts of Persuaficial dataset
In addition to the English results, we provide detailed F1 scores for persuasion detection on other language-specific subsets of the Persuaficial datasets. Table 11 reports results for German texts from SemEval 2023 Task 3 and their LLM-generated counterparts, Table 12 for French texts, Table 13 for Italian texts, Table 14 for Polish texts, and Table 15 for Russian texts. For each dataset, classifier performance on human-written texts (first column) is compared with performance on LLM-generated texts produced via paraphrasing, rewriting with subtle or intensive persuasion, and open-ended generation. The results are further broken down by both the generating model and the classifier model, demonstrating how different generation approaches influence the detectability of persuasion across languages.
Classifier Models Human-written Paraphrasing Generation Rewriting Generation Open-ended Generation
Subtle Persuasion Intensive Persuasion
Generating model: GPT 4.1 Mini
GPT 4.1 Mini 0.7203 $0.6999\downarrow$3% $0.4677\downarrow$35% $0.7959\uparrow$10% $0.9416\uparrow$31%
Llama 3.3 70B 0.7361 $0.7239\downarrow$2% $0.4574\downarrow$38% $0.8004\uparrow$9% $0.9347\uparrow$27%
[…]
Gemini 2.0 Flash 0.7903 $0.7768\downarrow$2% $0.6871\downarrow$13% $0.8244\uparrow$4% $0.8593\uparrow$9%
Generating model: Llama 3.3 70B
GPT 4.1 Mini 0.7203 $0.7102\downarrow$1% $0.4604\downarrow$36% $0.7836\uparrow$9% $0.9416\uparrow$31%
Llama 3.3 70B 0.7361 $0.7128\downarrow$3% $0.4681\downarrow$36% $0.7971\uparrow$8% $0.9347\uparrow$27%
[…]
Gemini 2.0 Flash 0.7903 $0.7816\downarrow$1% $0.7156\downarrow$9% $0.8298\uparrow$5% $0.8593\uparrow$9%
Generating model: Gemma 3 27b it
GPT 4.1 Mini 0.7203 $0.7303\uparrow$1% $0.3881\downarrow$46% $0.8996\uparrow$25% $0.9407\uparrow$31%
Llama 3.3 70B 0.7361 $0.7252\downarrow$1% $0.3779\downarrow$49% $0.8946\uparrow$22% $0.9339\uparrow$27%
Gemma 3 27b it 0.7655 $0.7877\uparrow$3% $0.6411\downarrow$16% $0.8824\uparrow$15% $0.8998\uparrow$18%
Gemini 2.0 Flash 0.7903 $0.7978\uparrow$1% $0.6644\downarrow$16% $0.8499\uparrow$8% $0.8584\uparrow$9%
Generating model: Gemini 2.0 Flash
GPT 4.1 Mini 0.7203 $0.7425\uparrow$3% $0.4476\downarrow$38% $0.9033\uparrow$25% $0.9416\uparrow$31%
[…]
Gemma 3 27b it 0.7655 $0.7866\uparrow$3% $0.6791\downarrow$11% $0.8806\uparrow$15% $0.9006\uparrow$18%
Gemini 2.0 Flash 0.7903 $0.7959\uparrow$1% $0.6949\downarrow$12% $0.8499\uparrow$8% $0.8593\uparrow$9%
Table 11: F1 scores for persuasion detection on German data sample of Persuaficial. More specifically, on sample of Persuaficial generated using German texts from SemEval 2023 Task 3 dataset. The first column reports performance on SemEval 2023 Task 3 German human-annotated texts. The remaining columns show performance on LLM-generated German counterparts.
Classifier Models Human-written Paraphrasing Generation Rewriting Generation Open-ended Generation
Subtle Persuasion Intensive Persuasion
Generating model: GPT 4.1 Mini
GPT 4.1 Mini 0.7505 $0.7356\downarrow$2% $0.4568\downarrow$39% $0.8021\uparrow$7% $0.9251\uparrow$23%
Llama 3.3 70B 0.7605 $0.7460\downarrow$2% $0.4675\downarrow$39% $0.7996\uparrow$5% $0.9164\uparrow$20%
Gemma 3 27b it 0.7827 $0.7733\downarrow$1% $0.6489\downarrow$17% $0.8190\uparrow$5% $0.8816\uparrow$13%
[…]
Generating model: Llama 3.3 70B
GPT 4.1 Mini 0.7505 $0.7160\downarrow$5% $0.4548\downarrow$39% $0.7959\uparrow$6% $0.9251\uparrow$23%
Llama 3.3 70B 0.7605 $0.7216\downarrow$5% $0.4794\downarrow$37% $0.7996\uparrow$5% $0.9174\uparrow$21%
Gemma 3 27b it 0.7827 $0.7744\downarrow$1% $0.6908\downarrow$12% $0.8343\uparrow$7% $0.8826\uparrow$13%
Gemini 2.0 Flash 0.7812 $0.7758\downarrow$1% $0.7116\downarrow$9% $0.8269\uparrow$6% $0.8418\uparrow$8%
Generating model: Gemma 3 27b it
[…]
Llama 3.3 70B 0.7605 $0.7366\downarrow$3% $0.3520\downarrow$54% $0.8815\uparrow$16% $0.9174\uparrow$21%
Gemma 3 27b it 0.7827 $0.8068\uparrow$3% $0.6179\downarrow$21% $0.8676\uparrow$11% $0.8826\uparrow$13%
Gemini 2.0 Flash 0.7812 $0.7898\uparrow$1% $0.6523\downarrow$17% $0.8388\uparrow$7% $0.8418\uparrow$8%
Generating model: Gemini 2.0 Flash
GPT 4.1 Mini 0.7505 $0.7807\uparrow$4% $0.4670\downarrow$38% $0.8996\uparrow$20% $0.9251\uparrow$23%
Llama 3.3 70B 0.7605 $0.7759\uparrow$2% $0.5120\downarrow$33% $0.8920\uparrow$17% $0.9174\uparrow$21%
Gemma 3 27b it 0.7827 $0.7920\uparrow$1% $0.6771\downarrow$13% $0.8696\uparrow$11% $0.8826\uparrow$13%
Gemini 2.0 Flash 0.7812 $0.7972\uparrow$2% $0.6858\downarrow$12% $0.8408\uparrow$8% $0.8418\uparrow$8%
Table 12: F1 scores for persuasion detection on French data sample of Persuaficial. More specifically, on sample of Persuaficial generated using French texts from SemEval 2023 Task 3 dataset. The first column reports performance on SemEval 2023 Task 3 French human-annotated texts. The remaining columns show performance on LLM-generated French counterparts.
Classifier Models Human-written Paraphrasing Generation Rewriting Generation Open-ended Generation
Subtle Persuasion Intensive Persuasion
Generating model: GPT 4.1 Mini
GPT 4.1 Mini 0.7471 $0.7511\uparrow$1% $0.4368\downarrow$42% $0.7959\uparrow$7% $0.9200\uparrow$23%
Llama 3.3 70B 0.7584 $0.7277\downarrow$4% $0.4447\downarrow$41% $0.7926\uparrow$5% $0.9166\uparrow$21%
Gemma 3 27b it 0.7659 $0.7775\uparrow$2% $0.6563\downarrow$14% $0.8185\uparrow$7% $0.8688\uparrow$13%
Gemini 2.0 Flash 0.7986 $0.7912\downarrow$1% $0.6750\downarrow$15% $0.8192\uparrow$3% $0.8410\uparrow$5%
Generating model: Llama 3.3 70B
GPT 4.1 Mini 0.7471 $0.7113\downarrow$5% $0.4812\downarrow$36% $0.7822\uparrow$5% $0.9200\uparrow$23%
Llama 3.3 70B 0.7584 $0.6823\downarrow$10% $0.4689\downarrow$38% $0.7827\uparrow$3% $0.9166\uparrow$21%
Gemma 3 27b it 0.7659 $0.7576\downarrow$1% $0.6967\downarrow$9% $0.8164\uparrow$7% $0.8688\uparrow$13%
Gemini 2.0 Flash 0.7986 $0.7827\downarrow$2% $0.7193\downarrow$10% $0.8192\uparrow$3% $0.8410\uparrow$5%
Generating model: Gemma 3 27b it
GPT 4.1 Mini 0.7471 $0.7239\downarrow$3% $0.3252\downarrow$56% $0.8925\uparrow$19% $0.9190\uparrow$23%
Llama 3.3 70B 0.7584 $0.7110\downarrow$6% $0.3376\downarrow$55% $0.8912\uparrow$18% $0.9156\uparrow$21%
Gemma 3 27b it 0.7659 $0.7944\uparrow$4% $0.6280\downarrow$18% $0.8619\uparrow$13% $0.8678\uparrow$13%
Gemini 2.0 Flash 0.7986 $0.7944\downarrow$1% $0.6356\downarrow$20% $0.8410\uparrow$5% $0.8401\uparrow$5%
Generating model: Gemini 2.0 Flash
[…]
Llama 3.3 70B 0.7584 $0.7479\downarrow$1% $0.4629\downarrow$39% $0.9015\uparrow$19% $0.9156\uparrow$21%
Gemma 3 27b it 0.7659 $0.7922\uparrow$3% $0.6632\downarrow$13% $0.8629\uparrow$13% $0.8688\uparrow$13%
Gemini 2.0 Flash 0.7986 $0.8069\uparrow$1% $0.6826\downarrow$15% $0.8410\uparrow$5% $0.8410\uparrow$5%
Table 13: F1 scores for persuasion detection on Italian data sample of Persuaficial. More specifically, on sample of Persuaficial generated using Italian texts from SemEval 2023 Task 3 dataset. The first column reports performance on SemEval 2023 Task 3 Italian human-annotated texts. The remaining columns show performance on LLM-generated Italian counterparts.
Classifier Models Human-written Paraphrasing Generation Rewriting Generation Open-ended Generation
Subtle Persuasion Intensive Persuasion
Generating model: GPT 4.1 Mini
GPT 4.1 Mini 0.7330 $0.7010\downarrow$4% $0.4634\downarrow$37% $0.7936\uparrow$8% $0.9372\uparrow$28%
Llama 3.3 70B 0.7676 $0.7398\downarrow$4% $0.5165\downarrow$33% $0.8016\uparrow$4% $0.9208\uparrow$20%
[…]
Generating model: Llama 3.3 70B
GPT 4.1 Mini 0.7330 $0.6829\downarrow$7% $0.5215\downarrow$29% $0.7793\uparrow$6% $0.9352\uparrow$28%
Llama 3.3 70B 0.7676 $0.7177\downarrow$7% $0.5637\downarrow$27% $0.7880\uparrow$3% $0.9198\uparrow$20%
Gemma 3 27b it 0.7728 $0.7596\downarrow$2% $0.7232\downarrow$6% $0.8176\uparrow$6% $0.8834\uparrow$14%
Gemini 2.0 Flash 0.7733 $0.7563\downarrow$2% $0.7321\downarrow$5% $0.7980\uparrow$3% $0.8217\uparrow$6%
Generating model: Gemma 3 27b it
GPT 4.1 Mini 0.7330 $0.7143\downarrow$3% $0.3777\downarrow$48% $0.9086\uparrow$24% $0.9372\uparrow$28%
Llama 3.3 70B 0.7676 $0.7330\downarrow$5% $0.4311\downarrow$44% $0.9047\uparrow$18% $0.9208\uparrow$20%
Gemma 3 27b it 0.7728 $0.7869\uparrow$2% $0.6737\downarrow$13% $0.8754\uparrow$13% $0.8834\uparrow$14%
[…]
Generating model: Gemini 2.0 Flash
GPT 4.1 Mini 0.7330 $0.7258\downarrow$1% $0.4696\downarrow$36% $0.9117\uparrow$24% $0.9372\uparrow$28%
[…]
Gemma 3 27b it 0.7728 $0.7916\uparrow$2% $0.6764\downarrow$12% $0.8704\uparrow$13% $0.8834\uparrow$14%
Gemini 2.0 Flash 0.7733 $0.7858\uparrow$2% $0.6975\downarrow$10% $0.8207\uparrow$6% $0.8217\uparrow$6%
Table 14: F1 scores for persuasion detection on Polish data sample of Persuaficial. More specifically, on sample of Persuaficial generated using Polish texts from SemEval 2023 Task 3 dataset. The first column reports performance on SemEval 2023 Task 3 Polish human-annotated texts. The remaining columns show performance on LLM-generated Polish counterparts.
Classifier Models Human-written Paraphrasing Generation Rewriting Generation Open-ended Generation
Subtle Persuasion Intensive Persuasion
Generating model: GPT 4.1 Mini
GPT 4.1 Mini 0.7246 $0.6889\downarrow$5% $0.4464\downarrow$38% $0.7844\uparrow$8% $0.9017\uparrow$24%
Llama 3.3 70B 0.7408 $0.7166\downarrow$3% $0.4556\downarrow$38% $0.7915\uparrow$7% $0.9071\uparrow$22%
[…]
Gemini 2.0 Flash 0.7683 $0.7571\downarrow$1% $0.6767\downarrow$12% $0.7784\uparrow$1% $0.8019\uparrow$4%
Generating model: Llama 3.3 70B
GPT 4.1 Mini 0.7246 $0.6889\downarrow$5% $0.4794\downarrow$34% $0.7546\uparrow$4% $0.9017\uparrow$24%
Llama 3.3 70B 0.7408 $0.6740\downarrow$9% $0.4635\downarrow$37% $0.7603\uparrow$3% $0.9091\uparrow$23%
Gemma 3 27b it 0.7360 $0.7203\downarrow$2% $0.6311\downarrow$14% $0.7732\uparrow$5% $0.8562\uparrow$16%
Gemini 2.0 Flash 0.7683 $0.7477\downarrow$3% $0.6986\downarrow$9% $0.7774\uparrow$1% $0.8019\uparrow$4%
Generating model: Gemma 3 27b it
GPT 4.1 Mini 0.7246 $0.7138\downarrow$1% $0.3651\downarrow$50% $0.8805\uparrow$22% $0.9017\uparrow$24%
Llama 3.3 70B 0.7408 $0.7248\downarrow$2% $0.3562\downarrow$52% $0.8858\uparrow$20% $0.9091\uparrow$23%
Gemma 3 27b it 0.7360 $0.7491\uparrow$2% $0.5907\downarrow$20% $0.8453\uparrow$15% $0.8562\uparrow$16%
[…]
Generating model: Gemini 2.0 Flash
GPT 4.1 Mini 0.7246 $0.7378\uparrow$2% $0.4660\downarrow$36% $0.8774\uparrow$21% $0.9017\uparrow$24%
Llama 3.3 70B 0.7408 $0.7500\uparrow$1% $0.4496\downarrow$39% $0.8920\uparrow$20% $0.9091\uparrow$23%
Gemma 3 27b it 0.7360 $0.7572\uparrow$3% $0.6227\downarrow$15% $0.8362\uparrow$14% $0.8562\uparrow$16%
[…]
Table 15: F1 scores for persuasion detection on Russian data sample of Persuaficial. More specifically, on sample of Persuaficial generated using Russian texts from SemEval 2023 Task 3 dataset. The first column reports performance on SemEval 2023 Task 3 Russian human-annotated texts. The remaining columns show performance on LLM-generated Russian counterparts.
Appendix H StyloMetrix and Its Usefulness in AI vs. Human Persuasive Text Analysis
To further demonstrate the utility of StyloMetrix for analyzing human-written versus AI-generated persuasive texts, we conduct a classification study using classical machine learning models with features calculated by StyloMetrix. We aimed to prove that linguistic features contain enough information to differentiate AI-generated persuasion from human-written persuasion.
For all GPT 4.1. Mini-generated English synthetic texts produced with each generation approach, we trained a separate classifier. For each experiemnt, we split the data into training and test sets, allocating 70% for training and 30% for testing. To ensure a credible evaluation, each human-written text and its LLM-generated counterpart were placed in the same split, either training or test. This prevents the classifier from exploiting the potential direct similarities between the paired texts. We employed widely used tree-based machine learning methods as classifiers, as they naturally capture non-linear interactions and are well-suited for moderate- to high-dimensional tabular data Grinsztajn et al. (2022); Uddin and Lu (2024). Previous work has shown that, for tabular data, tree-based models can even outperform deep learning approaches Grinsztajn et al. (2022).
Table 16 shows the results of these experiments. The outcomes show a clear progression: the more generative freedom the LLM is given, the easier it becomes for tree-ensemble models to distinguish its outputs from human-written persuasion. Paraphrasing keeps the AI text close to the original human style, yielding only moderate detection performance. In the Rewriting conditions, the model introduces larger stylistic shifts—whether by making persuasion subtler or more intense—which improves separability. Open-ended generation, starting from only a brief neutral summary, produces the greatest stylistic divergence and thus the highest classification accuracy. Overall, stylistic features become increasingly informative as the generation task becomes less constrained.
Model Precision Recall F1 Accuracy
GPT 4.1 Mini – Paraphrasing
RF 0.74 0.74 0.74 0.74
XGB 0.76 0.76 0.76 0.76
LGBM 0.76 0.76 0.76 0.76
GPT 4.1 Mini – Rewriting with Subtle Persuasive Effect
RF 0.84 0.84 0.84 0.84
XGB 0.87 0.87 0.87 0.87
LGBM 0.86 0.86 0.86 0.86
GPT 4.1 Mini – Rewriting with Intensified Persuasive Effect
RF 0.83 0.83 0.83 0.83
XGB 0.84 0.84 0.84 0.84
LGBM 0.86 0.85 0.85 0.85
GPT 4.1 Mini – Open-ended
RF 0.97 0.97 0.97 0.97
XGB 0.97 0.97 0.97 0.97
LGBM 0.98 0.98 0.98 0.98
Table 16: Classification performance of detecting GPT-4.1-Mini-generated persuasive texts versus human-written persuasive texts using linguistic-feature representations. Each experiment reflects a different AI-generation strategy and uses data combined from three English datasets.
Appendix I Cohen’s d statistic definition and our computation
For each linguistic feature, we calculate Cohen’s $d$ to quantify the magnitude of the shift between the feature distribution of generated texts $g_{i}$ and that of their human-written persuasive counterparts $r_{i}$. Cohen’s $d$ is defined in Equation 5 and is computed as follows.
First, we calculate the mean of each feature for human-written and generated texts:
$\bar{r}=\frac{1}{n_{r}}\sum_{i=1}^{n_{r}}r_{i},\qquad\bar{g}=\frac{1}{n_{g}}\sum_{i=1}^{n_{g}}g_{i},$ (1)
where $n_{r}$ and $n_{g}$ denote the number of human-written and generated texts, respectively. In our experiments, $n_{r}=n_{g}$ as for each human-written persuasive text we have AI-generated counterpart.
Next, we compute the sample standard deviations for each group:
$s_{r}=\sqrt{\frac{1}{n_{r}-1}\sum_{i=1}^{n_{r}}(r_{i}-\bar{r})^{2}},$ (2)
$s_{g}=\sqrt{\frac{1}{n_{g}-1}\sum_{i=1}^{n_{g}}(g_{i}-\bar{g})^{2}}.$ (3)
Using these, we calculate the pooled standard deviation:
$s_{\text{pooled}}=\sqrt{\frac{(n_{r}-1)s_{r}^{2}+(n_{g}-1)s_{g}^{2}}{n_{r}+n_{g}-2}}.$ (4)
Finally, Cohen’s $d$ is obtained as the difference in means normalized by the pooled standard deviation:
$d=\frac{\bar{g}-\bar{r}}{s_{\text{pooled}}}$ (5)
This effect size, Cohen’s $d$, provides a standardized measure of the shift in feature distributions between generated and human-written persuasive texts, allowing comparison across features and models.
Since many feature distributions may be non-Gaussian, we avoid this assumptions when testing for a significance of the shift. Statistical significance is assessed with a paired Wilcoxon signed-rank test on the per text differences $d_{i}=g_{i}-r_{i}$, with Benjamini-Hochberg FDR correction across features. As we perform one test per feature, we control for multiple comparisons using the Benjamini-Hochberg false discovery rate (FDR) correction Benjamini and Hochberg (1995) across all features. We report a significance indicator in our tables.
Appendix J Linguistic Differences between Human and Machine Generated Persuasive Texts - Additional Results
Tables 17 - 32 report the top features that most strongly distinguish AI-generated persuasive texts from human-written persuasive texts. We provide 16 tables in total, reflecting Cohen’s d effect sizes computed separately for each generating model and for each generation setting: Paraphrasing, Rewriting subtle persuasion, Rewriting intensified persuasion, and Open-ended generation.
Stylometric Feature Cohen’s d Sig.
GPT 4.1 Mini - Paraphrasing
L_CONT_T 0.5054
L_CONT_A 0.4590
SY_INV_PATTERNS -0.4542
LTOKEN_RATIO_LEM 0.3713
L_FUNC_A -0.3619
L_PUNCT_DASH 0.3378
ST_REPET_WORDS -0.3280
L_PUNCT_SEMC -0.2656
L_PUNCT_COM 0.2298
ASM -0.2196
VT_MIGHT 0.1995
L_LINKS -0.1954
L_ADJ_POSITIVE 0.1723
CDS -0.1719
PS_AGREEMENT -0.1667
L_ADV_SUPERLATIVE 0.1664
L_ADV_COMPARATIVE 0.1592
POS_ADV 0.1583
PS_CAUSE -0.1537
PS_TIME -0.1532
Table 17: Top 20 linguistic features for AI-generated persuasive text with Paraphrasing generation approach and GPT 4.1 Mini model vs. human-written persuasive texts (three samples of human datasets combined vs. AI counterparts). Cohen’s d sorted by absolute value.
Stylometric Feature Cohen’s d Sig.
GPT 4.1 Mini - Rewriting with Subtle Persuasive Effect
L_CONT_T 0.7036
L_CONT_A 0.6222
VT_MIGHT 0.6188
SY_INV_PATTERNS -0.6072
L_PLURAL_NOUNS 0.5444
L_FUNC_A -0.5293
SY_NARRATIVE 0.5242
LTOKEN_RATIO_LEM 0.4993
ST_REPET_WORDS -0.4319
POS_PRO -0.4307
L_YOU_PRON -0.4245
G_ACTIVE -0.4232
G_FUTURE -0.4015
L_SECOND_PERSON_PRON -0.3997
VT_FUTURE_SIMPLE -0.3949
CDS -0.3837
SENT_D_PP 0.3689
L_PUNCT -0.3401
VT_MAY 0.3378
SENT_ST_DIFFERENCE -0.3208
Table 18: Top 20 linguistic features for AI-generated persuasive text with Rewriting with Subtle Persuasive Effect generation approach and GPT 4.1 Mini model vs. human-written persuasive texts (three samples of human datasets combined vs. AI counterparts). Cohen’s d sorted by absolute value.
Stylometric Feature Cohen’s d Sig.
GPT 4.1 Mini - Rewriting with Intensified Persuasive Effect
L_CONT_T 0.7685
L_CONT_A 0.7654
L_PUNCT_DASH 0.7243
L_FUNC_A -0.5702
LTOKEN_RATIO_LEM 0.5075
L_ADV_SUPERLATIVE 0.5033
L_ADV_COMPARATIVE 0.4892
SY_INV_PATTERNS -0.4888
POS_ADV 0.4766
ST_REPET_WORDS -0.4037
L_ADJ_POSITIVE 0.3884
POS_ADJ 0.3670
ASM -0.3241
L_ADV_POSITIVE 0.3237
PS_CAUSE -0.3067
L_PUNCT_COM 0.3045
L_PUNCT_SEMC -0.2626
POS_PRO -0.2415
SENT_D_ADVP 0.2385
L_NOUN_PHRASES 0.2333
Table 19: Top 20 linguistic features for AI-generated persuasive text with Rewriting with Intensified Persuasive Effect generation approach and GPT 4.1 Mini model vs. human-written persuasive texts (three samples of human datasets combined vs. AI counterparts). Cohen’s d sorted by absolute value.
Stylometric Feature Cohen’s d Sig.
GPT 4.1 Mini - Open-ended
L_CONT_T 1.6204
L_CONT_A 1.4300
LTOKEN_RATIO_LEM 1.3172
SENT_D_NP 1.0532
ST_REPET_WORDS -0.9950
L_PUNCT_DASH 0.9885
L_FUNC_A -0.8585
SY_IMPERATIVE 0.8376
POS_NOUN 0.8076
POS_ADJ 0.7841
L_ADJ_POSITIVE 0.7816
SY_INV_PATTERNS -0.7496
L_LINKS -0.7474
VF_INFINITIVE 0.7252
G_PAST -0.6604
G_ACTIVE -0.6565
L_FIRST_PERSON_SING_PRON -0.6425
L_I_PRON -0.6425
SENT_D_VP 0.6275
VT_PAST_SIMPLE -0.5774
Table 20: Top 20 linguistic features for AI-generated persuasive text with Open-ended generation approach and GPT 4.1 Mini model vs. human-written persuasive texts (three samples of human datasets combined vs. AI counterparts). Cohen’s d sorted by absolute value.
Stylometric Feature Cohen’s d Sig.
Llama - Paraphrasing
SENT_ST_WPERSENT 0.7055
SENT_ST_DIFFERENCE -0.6979
L_PUNCT_COM 0.6659
SY_INV_PATTERNS -0.6146
L_CONT_T 0.5438
L_CONT_A 0.4861
G_ACTIVE -0.4821
L_PUNCT_DOT -0.4509
ASM -0.4192
POS_PRO -0.4055
FOS_FRONTING 0.3884
L_ADJ_POSITIVE 0.3844
L_PUNCT_SEMC -0.3839
L_YOU_PRON -0.3761
L_FUNC_A -0.3469
SY_SUBORD_SENT 0.3354
L_PUNCT -0.3339
POS_PREP 0.3256
VT_PRESENT_SIMPLE -0.3229
L_SECOND_PERSON_PRON -0.3228
Table 21: Top 20 linguistic features for AI-generated persuasive text with Paraphrasing generation approach and Llama 3.3 70B model vs. human-written persuasive texts (three samples of human datasets combined vs. AI counterparts). Cohen’s d sorted by absolute value.
Stylometric Feature Cohen’s d Sig.
Llama - Rewriting with Subtle Persuasive Effect
SY_INV_PATTERNS -0.7470
G_ACTIVE -0.7152
SENT_ST_WPERSENT 0.6961
L_CONT_T 0.6226
SENT_D_PP 0.6153
L_ADJ_POSITIVE 0.6145
POS_PREP 0.6086
L_YOU_PRON -0.5966
POS_NOUN 0.5896
L_PUNCT_COM 0.5848
L_SECOND_PERSON_PRON -0.5764
POS_ADJ 0.5595
SY_NARRATIVE 0.5419
L_CONT_A 0.5413
POS_PRO -0.5400
SENT_ST_DIFFERENCE -0.5382
SY_SUBORD_SENT 0.5373
L_PUNCT -0.5195
ASM -0.4923
L_PLURAL_NOUNS 0.4688
Table 22: Top 20 linguistic features for AI-generated persuasive text with Rewriting with Subtle Persuasive Effect generation approach and Llama 3.3 70B model vs. human-written persuasive texts (three samples of human datasets combined vs. AI counterparts). Cohen’s d sorted by absolute value.
Stylometric Feature Cohen’s d Sig.
Llama - Rewriting with Intensified Persuasive Effect
SENT_ST_WPERSENT 0.8019
L_ADJ_POSITIVE 0.7196
L_PUNCT_COM 0.7055
L_CONT_T 0.6589
POS_ADJ 0.6478
L_CONT_A 0.6349
G_ACTIVE -0.6171
SENT_ST_DIFFERENCE -0.6110
L_PUNCT_DOT -0.5876
SY_INV_PATTERNS -0.5135
ASM -0.4697
L_FUNC_A -0.4692
L_FUNC_T -0.4456
FOS_FRONTING 0.4390
POS_PRO -0.3973
SENT_D_VP 0.3895
L_PUNCT_SEMC -0.3806
CDS -0.3720
L_YOU_PRON -0.3670
L_NOUN_PHRASES 0.3630
Table 23: Top 20 linguistic features for AI-generated persuasive text with Rewriting with Intensified Persuasive Effect generation approach and Llama 3.3 70B model vs. human-written persuasive texts (three samples of human datasets combined vs. AI counterparts). Cohen’s d sorted by absolute value.
Stylometric Feature Cohen’s d Sig.
Llama - Open-ended
VF_INFINITIVE 1.3397
SY_IMPERATIVE 1.2406
G_ACTIVE -0.9149
SENT_D_VP 0.8331
L_PUNCT_DOT -0.8088
G_PAST -0.7696
L_OUR_PRON 0.7642
L_LINKS -0.7474
VT_PAST_SIMPLE -0.6985
SY_EXCLAMATION 0.6879
SY_INV_PATTERNS -0.6857
L_CONT_T 0.6780
SY_COORD_SENT 0.6774
L_PUNCT -0.6493
SENT_D_NP 0.6396
L_FIRST_PERSON_SING_PRON -0.6230
L_I_PRON -0.6230
L_WE_PRON 0.6190
L_IT_PRON 0.5548
VT_MUST 0.5078
Table 24: Top 20 linguistic features for AI-generated persuasive text with Open-ended generation approach and Llama 3.3 70B model vs. human-written persuasive texts (three samples of human datasets combined vs. AI counterparts). Cohen’s d sorted by absolute value.
Stylometric Feature Cohen’s d Sig.
Gemma - Paraphrasing
L_CONT_T 0.7526
L_CONT_A 0.6659
L_FUNC_A -0.6181
SY_INV_PATTERNS -0.6133
L_ADV_SUPERLATIVE 0.4274
ST_REPET_WORDS -0.4178
L_ADV_COMPARATIVE 0.4141
LTOKEN_RATIO_LEM 0.4081
PS_CONDITION -0.3995
ASM -0.3761
CDS -0.3606
POS_ADV 0.3605
L_ADJ_POSITIVE 0.3407
PS_AGREEMENT -0.3212
L_PUNCT_COM 0.3172
PS_CAUSE -0.2988
POS_ADJ 0.2872
POS_PRO -0.2805
G_ACTIVE -0.2710
SENT_ST_WPERSENT 0.2582
Table 25: Top 20 linguistic features for AI-generated persuasive text with Paraphrasing generation approach and Gemma 3 27b it model vs. human-written persuasive texts (three samples of human datasets combined vs. AI counterparts). Cohen’s d sorted by absolute value.
Stylometric Feature Cohen’s d Sig.
Gemma - Rewriting with Subtle Persuasive Effect
L_CONT_T 1.0245
L_CONT_A 1.0176
L_FUNC_A -0.8850
POS_PRO -0.7267
SY_INV_PATTERNS -0.7119
L_PLURAL_NOUNS 0.6762
POS_NOUN 0.6717
SENT_D_PP 0.6401
L_ADJ_POSITIVE 0.6191
G_ACTIVE -0.5837
L_SECOND_PERSON_PRON -0.5783
L_YOU_PRON -0.5736
SY_NARRATIVE 0.5666
POS_ADJ 0.5620
CDS -0.5438
L_FUNC_T -0.4852
VF_INFINITIVE -0.4846
ASM -0.4624
L_THEY_PRON -0.4486
POS_CONJ -0.4429
Table 26: Top 20 linguistic features for AI-generated persuasive text with Rewriting with Subtle Persuasive Effect generation approach and Gemma 3 27b it model vs. human-written persuasive texts (three samples of human datasets combined vs. AI counterparts). Cohen’s d sorted by absolute value.
Stylometric Feature Cohen’s d Sig.
Gemma - Rewriting with Intensified Persuasive Effect
L_CONT_T 1.0424
L_CONT_A 1.0122
L_FUNC_A -0.9718
L_ADJ_POSITIVE 0.9024
POS_ADJ 0.8267
PS_CONDITION -0.6942
PS_CAUSE -0.6071
L_FUNC_T -0.6021
SY_INV_PATTERNS -0.5923
G_ACTIVE -0.5882
L_NOUN_PHRASES 0.5486
POS_PRO -0.5389
L_ADV_SUPERLATIVE 0.5243
L_ADV_COMPARATIVE 0.5002
ASM -0.4781
CDS -0.4777
L_SINGULAR_NOUNS 0.4707
POS_NOUN 0.4570
L_YOU_PRON -0.4543
SENT_D_NP 0.4475
Table 27: Top 20 linguistic features for AI-generated persuasive text with Rewriting with Intensified Persuasive Effect generation approach and Gemma 3 27b it model vs. human-written persuasive texts (three samples of human datasets combined vs. AI counterparts). Cohen’s d sorted by absolute value.
Stylometric Feature Cohen’s d Sig.
Gemma - Open-ended
SY_IMPERATIVE 1.7354
L_CONT_T 1.4021
L_CONT_A 1.2523
SENT_D_NP 1.1601
PS_CAUSE -1.1529
VF_INFINITIVE 1.0451
L_FUNC_A -0.9623
PS_CONDITION -0.8951
POS_NOUN 0.8730
ST_REPET_WORDS -0.8055
LTOKEN_RATIO_LEM 0.7851
L_LINKS -0.7474
SY_INV_PATTERNS -0.7306
L_SINGULAR_NOUNS 0.6881
SY_EXCLAMATION 0.6825
SENT_D_VP 0.6631
POS_PREP -0.6571
L_I_PRON -0.6461
L_FIRST_PERSON_SING_PRON -0.6461
L_NOUN_PHRASES 0.6320
Table 28: Top 20 linguistic features for AI-generated persuasive text with Open-ended generation approach and Gemma 3 27b it model vs. human-written persuasive texts (three samples of human datasets combined vs. AI counterparts). Cohen’s d sorted by absolute value.
Stylometric Feature Cohen’s d Sig.
Gemini - Paraphrasing
L_CONT_A 0.7275
L_CONT_T 0.7074
SY_INV_PATTERNS -0.5802
L_FUNC_A -0.5371
L_PUNCT_COM 0.4880
LTOKEN_RATIO_LEM 0.4192
L_ADJ_POSITIVE 0.4104
ST_REPET_WORDS -0.3981
POS_ADJ 0.3677
L_NOUN_PHRASES 0.3255
ASM -0.2996
PS_CONDITION -0.2850
PS_CAUSE -0.2759
L_POSSESSIVES 0.2719
POS_PREP -0.2305
CDS -0.2242
L_PUNCT 0.2218
PS_AGREEMENT -0.2169
L_PLURAL_NOUNS 0.2053
L_THEIR_PRON 0.2050
Table 29: Top 20 linguistic features for AI-generated persuasive text with Paraphrasing generation approach and Gemini 2.0 Flash model vs. human-written persuasive texts (three samples of human datasets combined vs. AI counterparts). Cohen’s d sorted by absolute value.
Stylometric Feature Cohen’s d Sig.
Gemini - Rewriting with Subtle Persuasive Effect
L_CONT_A 0.9216
L_CONT_T 0.9163
L_FUNC_A -0.7257
SY_INV_PATTERNS -0.7182
POS_NOUN 0.5946
G_ACTIVE -0.5892
POS_PRO -0.5535
VT_MIGHT 0.5222
CDS -0.5034
SENT_D_PP 0.4976
L_YOU_PRON -0.4908
LTOKEN_RATIO_LEM 0.4813
L_PLURAL_NOUNS 0.4761
L_ADJ_POSITIVE 0.4695
L_SECOND_PERSON_PRON -0.4555
ASM -0.4517
L_PUNCT_COM 0.4347
ST_REPET_WORDS -0.4182
POS_ADJ 0.4165
SY_NARRATIVE 0.4079
Table 30: Top 20 linguistic features for AI-generated persuasive text with Rewriting with Subtle Persuasive Effect generation approach and Gemini 2.0 Flash model vs. human-written persuasive texts (three samples of human datasets combined vs. AI counterparts). Cohen’s d sorted by absolute value.
Stylometric Feature Cohen’s d Sig.
Gemini - Rewriting with Intensified Persuasive Effect
L_CONT_A 1.0170
L_CONT_T 0.9948
L_FUNC_A -0.8186
L_ADJ_POSITIVE 0.7827
POS_ADJ 0.7239
L_PUNCT_COM 0.5901
L_NOUN_PHRASES 0.5756
SY_INV_PATTERNS -0.5355
LTOKEN_RATIO_LEM 0.4892
L_ADV_SUPERLATIVE 0.4824
PS_CONDITION -0.4649
L_ADV_COMPARATIVE 0.4629
ASM -0.4428
ST_REPET_WORDS -0.4386
PS_CAUSE -0.4330
POS_ADV 0.4270
G_ACTIVE -0.4252
POS_PRO -0.3773
L_FUNC_T -0.3605
PS_AGREEMENT -0.3525
Table 31: Top 20 linguistic features for AI-generated persuasive text with Rewriting with Intensified Persuasive Effect generation approach and Gemini 2.0 Flash model vs. human-written persuasive texts (three samples of human datasets combined vs. AI counterparts). Cohen’s d sorted by absolute value.
Stylometric Feature Cohen’s d Sig.
Gemini - Open-ended
L_CONT_T 1.5669
L_CONT_A 1.3747
SY_IMPERATIVE 1.3186
LTOKEN_RATIO_LEM 1.1953
SY_EXCLAMATION 1.1092
VF_INFINITIVE 1.0846
ST_REPET_WORDS -1.0032
SENT_D_VP 0.9291
L_FUNC_A -0.8914
POS_NOUN 0.8779
L_NOUN_PHRASES 0.8190
PS_CAUSE -0.7878
PS_CONDITION -0.7832
G_ACTIVE -0.7769
L_LINKS -0.7474
L_SINGULAR_NOUNS 0.7331
SY_INV_PATTERNS -0.6693
G_PAST -0.6464
POS_PREP -0.6279
L_I_PRON -0.6236
Table 32: Top 20 linguistic features for AI-generated persuasive text with Open-ended generation approach and Gemini 2.0 Flash model vs. human-written persuasive texts (three samples of human datasets combined vs. AI counterparts). Cohen’s d sorted by absolute value.